Watch this page in the wiki to subscribe to automatic updates to this status page.
Please do not refer to any NCSA Industry Partners on this page. Please use the iforge nomenclature for all of the *forge infrastructure.
|START||END||What System/Service is affected||What is happening?||What will be affected?||Contact Person|
Upcoming Scheduled Maintenance
|Start||End||What System/Service is affected||What is happening?||What will be affected?||Contact Person|
|2019-01-23 0500||2019-01-25 |
|Scheduled Maintenance||ncsa#Nearline (GO) will be paused "display maintenance msg" Software and firmware email@example.com|
|Fileserver||Scheduled Maintenance||Shares on Fileserver will be unavailable during the outage. Documents & Data stored here will be firstname.lastname@example.org|
|2019-02-05 07:00||2019-02-05 19:00||iForge/cForge/SIMPL/aForge||Quarterly Maintenance||All systems will be unavailable during the email@example.com|
Previous Outages or Maintenance
|Start||End||What System/Service was affected?||What happened?||What was affected?||Outcome|
|2019-01-18 12:14||2019-01-18 14:32||RSA OTP user portal||An ESXi server crashed taking down several VMs it was hosting. The OTP VM rebooted on an alternate ESXi hosts.||RSA OTP user portal|
RSA OTP user portal online
|2019-01-18 12:14||2019-01-18 13:30||JIRA, file-server, ad-a, jabber, vsphere, email relay||An ESXi server crashed taking down several VMs it was hosting. The VMs all rebooted on alternate ESXi hosts.|
JIRA, file-server, ad-a, jabber, vsphere, and email relay all rebooted
JIRA had index files corrupted and took a while to repair those
JIRA, file-server, ad-a, jabber, vsphere, and email relay rebooted and online
|2019-01-17 08:00||2019-01-17 12:00||LSST|
|ALL LSST systems (incl. lsst-dev01, lsst-xfer, etc. as well as PDAC, verification, and Kubernetes clusters, and tus-ats01)|
Maintenance was completed successfully with the following caveats:
Please open tickets if you notice other issues.
|NPCF Emergency power off||Emergency power off panel was energized||Facility electrical and HVAC systems||Panel is firstname.lastname@example.org|
|BW/Mainframe resource||Hung threads on scratch/home, paused the scheduler, HSN requires full reboot to recover 9:30AM||Mainframe rebooted and all running jobs were lost.||BW returned to service 1PM||Timothy Bouvet|
|2019-01-10 5:55PM||code42 crashplan pro e services had update for dataloss bug with MS OneDrive||Code42 crashplan service was updatet to latest release to fix a dataloss problem with clients also running MS One Drive.||Backup services were interrupted for a few minutes while services updated||Now running Code42 email@example.com|
|2019-01-10 3:20PM||2019-01-10 5:00PM||DUO 2-Factor Auth||DUO Upstream vendor reported issues with their service.|
|NCSA systems that use DUO for 2FA||DUO brought their systems back firstname.lastname@example.org|
|2019-01-09 10:28 AM||2019-01-10 |
|BW/HPSS||Power event at NPCF and recovery from fallout||HPSS ncsa#Nearline||HPSS ncsa#Nearline RTS||Glasgow, James A|
|2019-01-09 10:28 AM||2019-01-09|
|BW/All Resources Down||Power event at NPCF and recovery from fallout||All BW Resources Down||Power Restored, All Resources Except HPSS RTSemail@example.com|
|Industry systems/ LSST systems||Power event at NPCF caused some Industry and some LSST systems to go offline||Running jobs on iforge and other systems||The affected systems have been returned to service and users are being notified of which jobs to rerun|
|NCSA office net firewall||Software upgrade on NCSA firewall and some config changes.||NCSAnet wireless, Wired network (closed and partially-closed nets). IllinoisNet wireless will remain available during the maintenance.||Firewall upgrade did not go through however all services have been restored. NetEng is investigating and will work with the vendor to figure out a firstname.lastname@example.org|
|code42 crashplan pro e services had update for security issues||Code42 crashplan service was updated with the latest security fixes||Backup services were interrupted for a few minutes while services updated||Now running Code42 email@example.com|
|OTP self-services site was down||Power on the hypervisor running the rsa otp self-service site was lost and the service didn't restart||PIN changes and new software distribution was unavailable||Now running updated version of software and all functionality was firstname.lastname@example.org|
|12/27/18 1:16PM||12/27/18 11:40PM||NPCF - 2 power blips (B transformers)||System has been returned to service.||Blue Waters ongoing jobs all terminated, scheduler paused while mainframe rebooted.|
After an absurdly long outage to perform a reboot, the system was returned to service. There were apparently issues on shutdown, and again on bringup with various hardware fallout.
|Blue Waters/ bwedge, bwds2 - rebooted on backup server|
bwdsm-dev - unresponsive?
|jack.internal.ncsa.edu stopped respondiing/crash, we’re having intermittent issues with ESXi hosts kernel dumping||VM server went down impacting vm's on that server. VM's will restart on other backup server with temporary interruption in their services.||jack.internal.ncsa.edu was power cycled. VM's were migrated to balance the load on servers after Jack was returned to service.|
|2018-12-17||2018-12-19||ICCP||Removed old file systems no longer in production; reformatted LUNs; rolling reboot of NSD servers to pick up new presentations; rebalance started||No user impact, all services remained fully operational||New v5 formatted disks added successfully; FS expanded to full size; rebalance of FS email@example.com|
|2018-12-18 08:00||2018-12-18 10:00||NPCF-EXIT-EAST||The firmware on NPCF-EXIT-EAST was upgraded.||Traffic was re-routed through NPCF-EXIT-WEST during the maintenance. No impact to users was observed.||Firmware was upgraded without firstname.lastname@example.org|
|2018-12-12 08:00||2018-12-12 20:43||ICCP|
Total cluster outage.
Taking a bit longer to bring the system back because interface renaming script stop working.
|2018-12-11 08:00||2018-12-11 10:00||NPCF-EXIT-WEST||The firmware on NPCF-EXIT-WEST was upgraded.||Traffic was re-routed through NPCF-EXIT-EAST during the maintenance. No user visible outage occurred.||Firmware was upgraded without email@example.com|
|2018-12-07 08:54||2018-12-07 19:25||ICCP||ACB UPS experienced fault causing storage appliance to shutdown in controlled manner||Jobs halted on system due to lack of parallel file system presence.||F&S was dispatched to fix put UPS in bypass, FSCK's were run on File Systems to ensure integrity and the cluster was returned to firstname.lastname@example.org|
|Wired network connections in NPCF office space||Software upgrade on network switches||wired network service in NPCF office space. NCSAnet, IllinoisNet Wireless remained available||Maintenance was email@example.com|
|2018-12-06 06:00||2018-12-06 07:30||NCSA ITS vSphere vCenter||ITS vSphere vCenter server was upgraded to latest version||All VMs remained online during the maintenance, but management through vCenter was unavailable from 06:18-07:25.|
Upgrade was completed successfully
|2018-11-29 08:00||2018-11-29 14:00||LSST|
ALL LSST systems (incl. lsst-dev01, lsst-xfer, etc. as well as PDAC, verification, and Kubernetes clusters, and tus-ats01)
|Maintenance was completed firstname.lastname@example.org|
|2018-11-19 08:00||2018-11-19 19:30||ICCP|
|Total cluster email@example.com|
|NCSA building router in 2045||software upgrade on one of the building routers (2045-br)||Traffic failed over to redundant building router and no impact on network traffic was seen||Maintenance was completed firstname.lastname@example.org|
|Blue Waters/Home filesystem||MDS issue||scheduler paused|
|Home file system RTS||Timothy Bouvet|
|idp.ncsa.illinois.edu||Upgrade Shibboleth IdP from v3.3.2 to v.3.4.1||ECP (command line) Duo authentication is now supported natively by Shib IdP software.||work completed a day early||Terrence Fleury|
2018-11-14 10:45 am
Blue Waters /Home filesystem
|Investigation ongoing- suspect HSN quiesce||/home, and new job starts during the scheduler pause||back in service at 11:20||Timothy Bouvet|
2018-11-06 12:25 pm
Networking NetSure DC Distribution System
Tape Library QBERT and DIGDUG
|De-energize distribution power panel DP-6C-020 to install new power panel PPC4|
Loss of power to the core network DC Distribution panel (B Side), the network is 2N power feed, no impact on the network due to redundancy.
Loss of power to two tape libraries, a temporary power feeds will be provided.
iForge system will be powered down for quarterly maintenance.
|work completed as expected||Mohammad Rantissi|
|Cluster reboot||Memory performance on most k8s nodes was in degraded state as a result of a power event that occurred over the weekend. Reseating the nodes in their chassis slots resolves the issue.||Systems rebooted and memory performance is back to email@example.com|
|2018-11-10 ~04:40||2018-11-10 ~04:45||iForge (select compute nodes)|
A power event caused some compute nodes to reboot
|Select skylake platform compute nodes, including 7 nodes in the skylake queue. Jobs running on those nodes would have been impacted.||Systems rebooted and brought themselves back firstname.lastname@example.org|
|2018-11-10 ~04:40||2018-11-10 ~04:45||LSST (lspdev and select L1 hosts)|
A power event caused some hosts to reboot:
lspdev/Kubernetes cluster was unavailable from ~04:40 until ~09:30
select L1 hosts rebooted
|Systems should be back online and functioning. Users are asked to create tickets if there are lingering email@example.com|
|NCSA building router in basement 07 (ncsa-07-br)||software upgrade on one of the building routers.||Traffic failed over to redundant NCSA building router. No impact on the network was observed||Maintenance was completed successfully without any firstname.lastname@example.org|
|2018-11-07 16:50||2018-11-07 17:00||NCSA Jira||Jira was rebooted to increase RAM.||NCSA's Jira was offline while it's RAM configuration is upgraded.||Upgrade was completed successfully without any email@example.com|
|2018-11-06 06:00||2018-11-06 21:45||iForge / aForge||Quarterly Maintenance (20181106 Maintenance for iForge)|
All systems were unavailable during the maintenance.
Maintenance was completed successfully:
NOTE: OFED was updated to v4 on the clusters during the PM. Some MPI software may need to be recompiled due to changes in libraries (e.g., libpsm_infinipath is no longer present in OFED v4). Frequently used openmpi installations have been updated to accommodate this change. Software compiled against affected MPI software may also need to be recompiled.
|2018-11-06 07:00||2018-11-06 09:00||NCSA VPN Service||The VPN was upgraded.||The NCSA VPN service was down for maintenance||The NCSA VPN has been firstname.lastname@example.org|
|wired networking on 4th floor in NCSA building||Software upgrade on network closet switches||Wired network, VOIP phones on 4th floor. NCSAnet Wireless remained available during maintenance window.||upgrade was completed successfully without any email@example.com|
|2018-11-02 3:30 AM|
GPFS issue. "ls /usr/local" hangs.
direct access to some directories under /iusr/local was OK.
ie. "ls /usrlocal/modules-3.2.9.iforge" was OK.
iforge login node is currently down.
New ssh connections are hanging.
There is the potential for issues with running jobs.
Scheduler has been paused.
Something odd going on with iforge020 was causing hangs.
Once iforge020 was rebooted, access to /usr/local was unlocked.
|2018-10-30 9:00 p.m.|
|NCSA DHCP||Patches||The DHCP server will be unavailable periodically for reboots and patching. Possible timeouts for DHCP, but generally no interruptions are firstname.lastname@example.org|
|wired networking on 3rd floor in NCSA building||software upgrade on network closet switches||wired network, VOIP phones. NCSAnet Wireless remained available during maintenance window.||code upgrade completed successfully without any email@example.com|
|2018-10-22 1:00pm||IDDS servers||Patches||XRAS admin/review/submit UIs, XDCDB Admin UI, NAPS||Patches firstname.lastname@example.org|
|2018-10-18 08:00||2018-10-18 12:00||LSST|
|maintenance completed email@example.com|
|2018-10-17 08:00||2018-10-17 18:00||ICCP|
|Total system outage||maintenance firstname.lastname@example.org|
|2018-10-17 15:40||2018-10-17 23:00||3rd Floor Networking||Portions of the third floor did not have network connectivity due to a switch malfunction.||Portions of the third floor are without network connectivity.||The issue has been email@example.com|
|Blue Waters||Maintenance to apply security Patches||All services for Blue Waters will be down except for ncsa#Nearline||Outage extended for 2 hours due to unexpected power loss to 3 rows of firstname.lastname@example.org|
|DUO 2-Factor Auth||DUO Upstream vendor has reported issues with their service.|
|NCSA systems that use DUO for 2FA might experience intermittent email@example.com|
|2018-10-15 7:30 am|
2018-10-15 11:00 pm
|Nebula, File-server||Power Loss in the NCSA building is causing issues with systems||Nebula web services are turned off, File-server is unavailable||Systems we brought back online and firstname.lastname@example.org|
|2018-10-15 07:35||2018-10-15 09:15||LSST||Power event -> host outage at NCSA 3003|
affected: all physical LSST hosts (and VMs) at NCSA 3003:
|2018-10-11 16:30||2018-10-11 17:00||crashplan backup service||crashplan was upgraded to code42 6.8.4||crashplan service was restarted and clients reconnected||crashplan service has fewer security vulnerabilities email@example.com|
|2018-10-09||2018-10-09||DHCP||Additional DHCP attributes will be passed to clients.||The Security Operations group has requested that the Web Proxy Auto-Discovery Protocol (WPAD) be set to blank via DHCP to better secure client workstations/laptops. This should not impact any users general network usage.|
WPAD has been applied to all user networks at NCSA and NPCF (including wireless).
|2018-10-08 17:00||2018-10-08 21:00||Wired networking on 2nd floor in NCSA building||ncsa-2045 Network switch software upgrade||Wired networking for desktop computers and VOIP phones. Wireless network remained available during maintenance||switch stack on second floor was upgraded. There were some issues during upgrade process due to which maintenance ran longer than expected. All networking services are restored back to firstname.lastname@example.org|
|2018-10-4-16:35||2018-10-4-16:35||jabber.ncsa.illinois.edu||The open fire jabber server stopped working correctly and was restarted.||Everyone using jabber reconnected.||Jabber rooms are working like they should email@example.com|
|2018-10-04 08:00||2018-10-04 09:15||LSST||Critical security patching|
ALL LSST systems (incl. lsst-dev01, lsst-xfer, etc. as well as PDAC, verification, and Kubernetes clusters)
The following systems will remain online and unaffected:
|Maintenance was firstname.lastname@example.org|
|2018-10-03 06:00||2018-10-03 07:00||Campus Cluster - Networking||Maintenance was performed on the OmniPoP uplink on ur1carne, which is the upstream router for all ICCP based network traffic. Engineers worked to transition the link from old optical transport gear to new gear that is optically protected with automatic failover.||All traffic that would normally take this OmniPoP link will reroute through other WAN links on ur1carne. Downtime of < 15 min is expected within the hour window while engineers swing the fiber jumpers from the old optical gear to the new optical gear. There should be no impact to DES or any ICCP customers. Please contact NetEng if you notice any unexpected outages.||Maintenance was email@example.com|
|2018-10-02 17:00||2018-06-02 20:00||NPCF Networking DC Power System||Testing and maintenance of the DC power system and battery backup will be performed.||No outage.||Tests were completed without firstname.lastname@example.org|
|2018-09-26 11:00||2018-09-26 12:00||Campus Cluster - MWT2 Networking||Maintenance was performed on the Internet2 uplink on ur1carne, which is the upstream router for all ICCP/MWT2 based network traffic.||MWT2 lost connectivity to LHC1 but everything else rerouted, all of which was expected.||The maintenance was successful, no issues have been email@example.com|
|2018-09-20||2018-09-24||OpenAFS servers||OpenAFS file and database servers were upgraded to 1.6.23||The OpenAFS servers were upgraded to the latest code without service interuption||Now running with latest security fixes in firstname.lastname@example.org|
|2018-09-20 08:00||2018-09-22 16:50||LSST Qserv||qserv-master01 is having trouble booting after a motherboard replacement during planned maintenance.||Qserv in general, specifically qserv-master|
|2018-09-20 08:00||2018-09-20 14:40||LSST LSPdev|
LSPdev kubernetes is having a gateway error after upgrading
|2018-09-20 08:00||2018-09-20 14:00||LSST|
Monthly maintenance (Sep):
|All LSST systems and services will be unavailable for the duration of the maintenance period.|
qserv-master01 and LSPdev are still having issues. These will be tracked as a separate incidents.
|2018-09-19 08:00||2018-09-19 22:00||Campus Cluster|
All compute and login nodes were down.
The filesystems were also unavailable due to issues with the change to gpfs and RH7.5
|The cluster was back in service at email@example.com|
|2018-09-17 17:30||20018-09-19:30||Wired networking on 1st floor (ncsa-1045)||software upgrade on network switch for 1st floor.||Wired networking for users on 1st floor was unavailable as network engineering performed software upgrades on their equipment.|
Wireless network (NCSAnet) remained available during this time.
|Maintenance was completed successfully. Users can contact neteng if they have any issues with their wired network firstname.lastname@example.org|
|2018-09-12 06:00||2018-09-12 09:00||DNS1, DNS2||DNS1 and DNS2 will be updated/upgraded|
DNS servers will be undergoing routine maintenance. During this maintenance window, system
and services will be restarted. One DNS server will always be responsive during the maintenance.
|Updates have been email@example.com|
|2018-09-11 9:30 a.m.||2018-09-11 11:00 a.m.||Internet2 100G connection||ICCN engineers will be migrating our Internet2 connection to the new ICCN optical equipment.||Traffic will fail over to a secondary peering. We expect minimal impact to users. Direct peering will fall back to normal routing.||The migration has been firstname.lastname@example.org|
|2018-09-11 8:30 a.m.||2018-09-11 11:00 a.m.||ESnet 100G direct connection||We will be migrating our ESnet connection to the new ICCN optical equipment.||Traffic will fail over to a secondary peering. We expect no impact during this maintenance.||The migration has been completed.|
|netact.ncsa.illinois.edu||Multiple users reported they were unable to delete their activations or change networks within Netact.||netact.ncsa.illinois.edu||Fixed the bug and tested. Issue was email@example.com|
|2018-09-06 11:00||2018-09-06 12:00||MREN Circuit Move||The MREN WAN circuit is being moved to an optical protection switch.||Traffic will be re-routed over an alternate peering during the test firstname.lastname@example.org|
|2018-09-06 16:00||2018-09-06 16:40||RSA Authentication Manager||RSA Authentication Manager 8.2 SP 1 P 08 was applied||Both primary and replica servers were updated with the latest security patches||Running 8.2 SP1 P08email@example.com|
|2018-08-15 08:00||2018-08-15 20:08||Campus Cluster|
Corrected bad inode on filesystem.
Rebooted IB core switch
2x100G links are working
|2018-08-29 09:38||2018-08-29 10:21||Services that utilize Duo 2FA including bastions hosts and VPN.||latency issues with DUO1 as per https://status.duo.com/||any service that uses Duo for authentication including bastion hosts and VPN.||Service appears be to be returning to normal as per updates on https://firstname.lastname@example.org|
|2018-08-20 10:20||2018-08-20 12:00||sslvpn.ncsa.illinois.edu||intermittent login issues with DUO two factor authentication due to an outage on DUO's end.||Two factor authentication to sslvpn service.||Duo identified the issue and resolved the outage. Users can connect to sslvpn over Duo 2FA email@example.com|
|2018-08-16 11:25||2018-08-16 12:41||Slack||Slack is reporting connectivity issues on their status page (https://status.slack.com/)||Slack reported, "connectivity issues impacting all workspaces"|
Slack reported this resolved at 12:41, though NCSA users reported it working around 11:38.
|2018-08-15 08:00||2018-08-15 16:00||ISDA VM infrastructure||Upgrade of all VM servers as well as backend storage system||NCSA opensource, NCSA docker hub, ISDA VM servers||Upgrade was firstname.lastname@example.org|
|2018-08-15 0800||2018-08-15 1430||Storage Condo Maintenance||All servers were upgraded to gpfs 18.104.22.168 and the clustered nfs service was implemented as well.||Storage Condo||Upgrade was email@example.com|
|2018-08-14 05:00||2018-08-14 09:00||NCSA Wiki||wiki.ncsa.illinois.edu will be upgraded to Confluence 6.10.1 and then to 6.10.2.||The wiki will be down intermittently during the upgrade. Read the banner at the top of wiki pages for current status.||Upgrade was firstname.lastname@example.org|
|2018-08-07 07:00||2018-08-10 12:00||iForge ifdbpoc server||Hardware issues require migrating to new server; some signs indicate service was impacted prior to 2018-08-07 07:00 but no reports have confirmed||ifdbpoc||Admins migrated data and services to another server. Verification was performed by the apps team.|
|2018-08-08 -- 1430hrs||2018-08-10 -- 0730hrs||Blue Waters Nearline Endpoint|
Due to very high demand for data retrieval from Nearline, a pause rule is in effect to allow manual task scheduling. You may submit tasks as normal and they will be run as quickly as possible.
|Data storing and retrieving to/from the Nearline storage system.||Many tasks were manually scheduled and completed to help re-balance the system utilization. The endpoint pause rule was lifted and all tasks are running email@example.com|
|2018-08-07 07:00||2018-08-07 22:15||iForge / aForge||Quarterly Maintenance (20180807 Maintenance for iForge)||All systems will be unavailable during the maintenance.|
|2018-08-03 11:30||2018-08-03 13:30||NCSA VPN||A configuration issue caused some VPN users connection problems to some NCSA resources.||Some VPN users reported connectivity problems to some internal NCSA resources.||A configuration change was applied which corrected the routing firstname.lastname@example.org|
|2018-07-27 11:45||2018-07-27 13:45||NCSA Wiki||The wiki was being intermittently slow and unresponsive.||wiki.ncsa.illinois.edu||Upgraded several software packages and rebooted wiki email@example.com|
|2018-07-27 08:00||2018-07-27 08:15||NCSA VPN||The old NCSA VPN (vpn.ncsa.illinois.edu) was decommissioned. All users should be using the new VPN (sslvpn.ncsa.illinois.edu).||VPN was decommissioned.||The old VPN has been decommissioned and all users should be using the new VPN.||firstname.lastname@example.org|
|2018-07-26 14:00||2018-07-26 19:00||NCSA RT||The RT help site was being intermittently slow and unresponsive.||help.ncsa.illinois.edu||Upgraded several software packages and rebooted RT email@example.com|
|2018-07-25 14:30||2018-07-25 14:40||NCSA Wiki||Wiki Restart||Confluence service firstname.lastname@example.org|
|2018-07-24 13:20||2018-07-24 13:55||crashplan||crashplan was upgraded to 6.7.3 for latest feature and security updates. Client updates will push out to system automatically over the next few days.||all client paused backups for about 2 mins as servers restarted with new code.||now running Code42 email@example.com|
|2018-07-22 19:14||2018-07-22 19:45||NCSA GitLab||NCSA GitLab server was updated.||Completedfirstname.lastname@example.org|
|2018-07-19 18:44||2018-07-20 10:45:13||nebula||nebula controller experienced a fatal hardware error on 10gE nic|
horizon interface to nebula https://nebula.ncsa.illinois.edu/ and all open stack command line tools are non-functional. Keystone authentication services are also off-line.
Instances that were running should continue to run but restarting will probably fail until the controller is repaired. launching new instances will also fail.
|Replaced card, nebula.ncsa.illinois.edu is now accessible email@example.com|
|2018-07-19 12:00||2018-07-19 12:30|
LSST: lsst-dev-db and dependent services, including kubernetes lspdev
Following the July 19 planned maintenance, MariaDB services on lsst-dev-db are unavailable along with dependent services, including:
DB services on lsst-dev-db along with dependent services, including:
|2018-07-19 08:00||2018-07-19 12:00||LSST|
Monthly maintenance (July):
ALL lsst-dev systems (incl. lsst-dev01, lsst-xfer, etc. as well as PDAC, verification, and Kubernetes clusters)
The following systems will remain online and unaffected:
Maintenance was successfully completed, although the following resultant issue is being tracked in a separate status event:
DB services on lsst-dev-db are unavailable along with dependent services, including:
|2018-07-16– 900||2018-07-16– 1938||Bluewaters||System was upgraded for security issues and to migrate to Cuda 9.1||Bluewaters compute and scheduler||Bluewaters is now updated|
|2018-07-09 – 1130||2018-07-10 – 1700||Campus Cluster Monitoring Webpage||SET is moving set-analytics to https. This should have been a simple change to a host name, but after the change the new value was not picked up.||The monitoring web page gave a loading circle that never resolved to anything.||Set up a Grafana instance for the display of the Campus Cluster firstname.lastname@example.org|
|2018-06-28||2018-07-09||Nebula||Nebula was taken offline to repair the filesystem||All Nebula services||Nebula is performing well email@example.com|
|2018-06-29 -- 1300hrs||2018-07-08 – 1400hrs||Blue Waters Nearline Endpoint|
Due to very high demand for data retrieval from Nearline, a pause rule is in effect to allow manual task scheduling. You may submit tasks as normal and they will be run as quickly as possible.
|Tasks submitted to Globus will start in a paused state but will be released to run, at the earliest possible time, based on resource availability.||Backlog of file stages was cleared and endpoint pause rule firstname.lastname@example.org|
|Access to NPCF||For the July 4th UIUC fireworks show, the parking Lots E14 and E14-shuttle will be closed from 6:00 p.m. Monday, July 2nd, through 6:00 a.m. Friday, July 6th. No parking will be allowed in these locations at any time during this period. Please do not park in the NPCF dock area - use the shuttle buses, or park in lot E46 (south on Oak St.).||Parking facilites for NPCF||Parking is back to normal|
|2018-05-03 14:30||2018-06-28 09:00||iForge gpu queue||both nodes in the general 'gpu' queue were offline due to issues with the GPUs||iForge 'gpu' queue could not be used||Tried driver updates and engaged with vendors; ultimately got one node working with 4 M40 GPUs rather than the previous 2 K80 GPUs; continue engaging with vendors to get the other node working but queue is now available.|
|0800 2018-07-02||1200 2018-07-02||Blue Waters Nearline||One tape library (of four) will be powered down for hardware maintenance (replacement of tape import/export module).||Access to tapes in the affected library will be blocked until the system returns to service. Users staging data may see delays in accessing data until the library is back online.||Work was completed with some delay (scheduled to complete by 0930) due to a failed SD card (used for storing and loading library geometry).||email@example.com|
|2018-06-27 9:00||2018-06-27 1:00||LSST - k8s lspdev||kub001 unplanned reboot and kub004 ran out of memory.||lspdev JupyterHub|
Kubernetes pods restarted.
|2018-06-27 08:30||2018-06-27 11:49||Slack||Slack is reporting connectivity issues on their status page (https://status.slack.com/)||Slack||Slack reports, "workspaces should be able to connect again"||firstname.lastname@example.org|
|Blue Waters Scratch Filesystem||Top of Rack network switch died in rack 8. Cray onsite and performed a work around and will replace Monday. Sonexion rack 28 lost mind and was rebooted.||Partial scratch outage of ost169-179||bypassed faulty switch, rack 28 sonexion rebooted. faulty swich replaced Monday email@example.com|
|2018-06-21 -- 1200hrs||2018-06-23 -- 1045hrs||Blue Waters Nearline Endpoint|
Due to very high demand for data retrieval from Nearline, a pause rule is in effect to allow manual task scheduling. You may submit tasks as normal and they will be run as quickly as possible.
|Tasks submitted to Globus will start in a paused state but will be released to run, at the earliest possible time, based on resource availability.||Many tasks were pushed through the system by manually ordering them to reduce tape drive competition. Endpoint pause rule removed and all tasks resumed.||hpssadmin@ncsa.Illinois.edu|
|2018-06-21 08:00||2018-06-21 09:35||LSST|
Monthly maintenance (June):
CentOS 6.9 servers:
Other impact was not expected but unexpected issues could have lead to connectivity issues for other hosts or downtime for lsst-dev01 or hosted VMs
|Maintenance was firstname.lastname@example.org|
|2018-06-20 14:00||2018-06-20 19:00||Campus Cluster||Rolling reboot of the core IO servers to move GPFS from 22.214.171.124 to 126.96.36.199 for CentOS 7.5 support; No downtime occurred||Successful Upgrade||Cluster now supports CentOS 7.5 email@example.com|
|2018-06-18||2019-06-20 7pm||Nebula||Nebula was shut down to fix broken filesystems.||All Nebula services||Nebula is up and running again. Please contact firstname.lastname@example.org if you still see email@example.com|
|2018-06-19 08:00||2018-06-19 12:00||LSST L1 Test Stand|
Level One Test Stand, including:
|Maintenance completed firstname.lastname@example.org|
|2018-06-18 07:00||2018-06-18 09:30||vSphere & Various VMs||Two of our hosts went down with network interface errors.||Multiple VMs hosted on those nodes (incl. Fileserver, ncsa-print, and subversion)||Both hosts are back online as well as all VMsemail@example.com|
|2017-06-16 22:18:32||2017-06-17 08:10:00|
PBSPro server was hung on cfsched
|Job scheduling and job submission were failing.||restarted PBSPro server on cfsched||Jim Long|
|2018-06-15 1330hrs||2018-06-15 1530hrs||Blue Waters Nearline||Replacement of a tape robot transporter||This work is not expected to impact operations. The library system will continue to operate with a single transporter but mount times may be somewhat longer until the second unit is returned to firstname.lastname@example.org|
|2018-06-12 04:30||10:00||Blue Waters||Thunderstorms have resulted in a power interruption. This outage impacts both the compute nodes and all filesystems. Therefore, a full reboot will be necessary.Return to service is estimated to be approximately 10 am Central time.||Blue Water in total||Full reboot|
|2018-06-12 ~03:45||2018-06-12 ~06:00||Campus Cluster||Many compute nodes rebooted. No system on UPS was affected, and some compute nodes remained up. Facilities at ACB report that there were no power events this morning or last night, but this seems the most likely cause.||Many compute nodes, but not all. Jobs on the nodes that rebooted were lost.||Nodes rebooted at a similar time, and many returned in a state unsuitable to run jobs. Rebooting in smaller groups got everything working email@example.com|
|2018-06-12 ~03:40||2018-06-12 ~06:30||iForge|
A storm caused a brief power event which impacted:
|All nodes in the big_mem and skylake queues were rebooted by the power event.||Nodes rebooted on their own and were marked back online in the scheduler by around ~6:30am.|
|2018-06-12 ~03:40||2018-06-12 09:00||LSST|
Storm caused power event which impacted:
The following nodes rebooted because of the power event:
|2018-06-11 08:30||2018-06-11 8:35||Campus Cluster ADS||Vlan changes on campus cluster||campus cluster - Active data storage (ADS)||Maintenance completed firstname.lastname@example.org|
|2018-06-07 06:30||2018-06-07 14:00||Blue Waters||The boot node crashed requiring the system to be rebooted. File system and ESLogins remain up.||All running jobs were lost, no new jobs were started until system is return to service, Torque was updated to ver. email@example.com|
|2018-06-01 00:50||2018-06-01 03:50||Blue Waters||/var space filled up by additional logging in Moab to troubleshoot job slide issue.||PBS server went down due to no space in /var||Zipped and moved old Moab logs to lustre file system to free up /var space, then restarted PBS firstname.lastname@example.org|
|2018-05-31 14:00||2018-05-31 14:10||NCSA Open Source||Retirement of both HipChat and FishEye/Crucible||Services will be shutdown and archived.||Services are disabled and will be archived in a month.|
|2018-05-31 11:55||NCSA ITS vSphere vCenter||ITS vSphere vCenter server will be upgraded to the latest VMware vCenter 6.7||All VMs will remain online during the maintenance, but management through vCenter will be offline during the upgrade.||Successful upgrade to VMware vCenter email@example.com|
|2018-05-23 06:55||2018-05-24, 1900hrs||Campus Cluster File System||A failure of both disk array controllers serving the CC file systems resulted in abrupt loss of access to the underlying storage. One array controller was identified as broken while the storage system was brought back up on the remaining controller for inspection and analysis. A thorough check of the file systems and storage devices was started. At 1100hrs May 24th the replacement array controller arrived and was installed. After further testing to assure system stability, the file systems were brought back online and released to the cluster admins.||All campus cluster file systems||Normal cluster operations were resumed. Investigation into the root cause is ongoing with the cooperation of the system manufacturer.|
|2018-05-21||DNS1/2||There were a few reports of intermittent DNS lookups failures/slowness||Firewall state tables resources were being exhausted. Limits for those state tables have been increased. This appears to have resolved the problem.||No further reports of the issue, after making the firstname.lastname@example.org|
|ifsm.ncsa.illinois.edu System is being upgraded and rebooted||No services should be affected||yum upgrade and reboot|
|2018-05-17 8:00||2018-05-17 15:00||NPCF-Core-East||The hardware and firmware on the core east router was be upgraded||Traffic rerouted through npcf-core-west during the maintenance window. There was an unexpected outage for about 10 mins which impacted network connectivity throughout NCSA.||Upgrade on core-east was completed successfully. No further network outages are expected.|
|2018-05-09 7:00||2018-05-09 17:40||dns1.ncsa.illinois.edu||Enabling BIND on ipv6 and enabling a firewall on the server||No impact is expected.||Maintenance was completed.|
Monthly maintenance (May):
|All systems (except lsst-daq, lsst-l1-*, & tus-ats01) were unavailable for maintenance.|
Maintenance was extended until 13:30 and then completed.
External Grafana monitoring (monitor-ncsa.lsst.org) was offline until 14:25 due to storage rebuild on lsst-monitor01.
|2018-05-17 10:13||2018-05-17 10:18||Core Outage||During core router maintenance the incorrect core router was powered off.||Network connectivity across NCSA was affected.||The core router was powered back on, verified and brought back into service.|
|2018-05-16 17:40||Campus Cluster|
Monthly maintenance (May)
|Entire system was unavailable for maintenance.||Maintenance complete, all tasks complete.|
|2018-05-16 1100hrs||2018-05-16 1300hrs||ADS||Planned Campus Cluster network upgrades also impacted access to ADS||All ADS storage exports became unreachable||Eric has notified us that the networking maintenance is complete and ADS customers are able to access their storage again.|
|21 Mar 2018||14 May 2018||openxdmod.ncsa.illinois.edu||An update to Torque broke the updates of XDMoD. openxdmod.ncsa.illinois.edu was offline while the system it resided on was updated, all the dependency software was installed, and the latest version of XDMoD was installed. Then all the data had to be re-imported.||Software updating||Service restored with updated software.|
|2018-05-08 0000||2018-05-09 0015||NCSA Storage Condo|
One node ran out of memory, causing a deadlock in GPFS. During deadlock recovery, GPFS shut down on multiple nodes. Upon restart of the cluster, a different metadata server had a check on its PCI bus, forcing another unmount. All file systems but one were recovered. While recovering the last one, one of the Roger NetApp storage arrays started throwing errors, requiring a power cycle of the controller and disks, prompting a final recovery of the last file system.
|Condo file systems and services.||All file systems recovered and services restored.|
|2018-05-08 07:00||2018-05-08 07:40||iForge||Quarterly Maintenance (20180508 Maintenance for iForge)||All systems were unavailable during the maintenance.||Planned maintenance completed successfully|
|2018-05-08 8:00||2018-05-09 8:00||NPCF-Core-West||The hardware and firmware on the core router will be upgraded||Traffic will be rerouted through npcf-core-east during the maintenance window. No impact is expected.||The hardware and firmware was upgraded on npcf-core-west without incident. Traffic has been successfully failed back.|
|2018-05-03 08:45||2018-05-03 10:15||NCSA WIKI, JIRA, services that rely on NCSA LDAP||Large amount of connections from two particular servers were hitting LDAP, causing the slow-down that in term caused timeouts for various applications using LDAP authentication. Blocking the cuplrit servers remedied the situation||NCSA WIKI, NCSA JIRA, other applications that rely on NCSA LDAP authentication.||Culprit servers were blocked|
|9:00am||9:25am||syslog-sec.ncsa.illinois.edu||out of cycle patching of Security Syslog collectors to address CVE-2018-1000140||Load balance fail over to secondary collector, RELP will be buffered.|
relay-01 was updated and loadbalancer failed back.
|4/25 14:00||4/25 15:00||MREN WAN Circuit||WAN circuit testing.||Traffic will be re-routed over an alternate peering during the test period.||The MREN circuit was brought back in to production.|
|2018-04-24 12:30||2018-04-24 16:00||NCSA jabber service||jabber was down while we repaired its authorization configuration.||jabber.ncsa.illinois.edu wasn't accepting jabber logins||jabber working again.|
|2018-04-24 09:10||2018-04-24 09:50||LSST||increased LDAP timeout to 60 seconds in sssd.conf to fix problems with long login times and failure to start batch jobs||kub*, verify-worker*|
sssd.conf updated, sssd restarted
verify-worker nodes were drained during the change
affected nodes may have slow LDAP response times for a short while (due to local cache needing rebuilt)
|04/18/2018 10:30||04/18/2018 11:30||ICCP April Maintenance||Replaced 4x10G links from cc-core0 to carne. Updated BIOS on remaining parts of Cluster nodes.||No outage.||Completed without any outage.|
|04/18/2018 10:30||04/18/2018 11:30||ICCP core switches||One of the 4x10G links from cc-core0 to carne had incrementing errors and has been administratively down to prevent those errors from affecting traffic. There was a scratched fiber that earlier diagnosis had revealed, so we replaced the fiber during this ICCP PM.||Nothing, all traffic rerouted through cc-core1||The errors are still incrementing, but we've narrowed down the remaining options for what might be going on.|
|4/12 0930||4/12 1830||ADS NFS/Samba||The ESXi Hypervisor server had an error on it: 'A PCI error requiring a reboot has occurred.'.||ADS NFS/Samba/Gridftp||The server was rebooted, the error cleared and all systems/services were restarted.|
|4/11 03:00 p.m.||4/11 03.15 pm||Netact||Netact code was updated. Going forward new office activation names will have "-ofc" appended to them.||No service impact to Netact.||Change was successfully implemented. Netact remained in service during and after the change.|
|4/11 9:00||4/11 10:00||LSST NPCF Firewall||Primary firewall will be upgraded to use FRR instead of openBGP.||No impact is expected. The firewalls do not need to be failed over and no interruption in traffic flow is anticipated.||Firewall was successfully migrated. No downtime occurred.|
|dns1.ncsa.illinois.edu||OS Patching and BIND updates||dns1 (secondary DNS server) will be rebooted to apply patches. DNS2 will remain up.||DNS1 OS patching is completed. BIND was upgraded to 9.11. BIND is only bound currently to its ipv4 interface.|
|dns2.ncsa.illinois.edu||OS Patching and BIND updates||dns2 (secondary DNS server) will be rebooted to apply patches. DNS1 will remain up. An IPv6 address will also be added to system in preparation for a broader IPv6 DNS rollout.||DNS2 OS was patched. BIND was upgrade to 9.11. IPv6 Address was also enabled on the server and BIND is listening on that address.|
|MREN WAN Circuit||Port Move||Traffic will be re-routed over an alternate peering during the maintenance.||The port was moved and the circuit was brought back into service without issue.|
|LDAP||LDAP process crashed||Authentication to LDAP-backed services||LDAP was upgraded and restarted|
|MREN WAN Circuit||Port Move||Traffic will be re-routed over an alternate peering during the maintenance.||The port was moved and the circuit was brought back into service without issue.|
|MREN WAN Circuit||WAN circuit testing.||Traffic will be re-routed over an alternate peering during the test period.||Testing was completed and the circuit was brought back into service.|
|2018-03-21 08:00||2018-03-21 17:30||Campus Cluster manage server and compute nodes except DES and MWT2||Deploying new management server, upgrading to Torque 6.1.2 and Moab 9.1.2. Bios update. Configuration changes on GPFS servers. Tech Service CARNE code upgrade.||Scheduler down. User access disabled|
New management server is up with Centos7. Installed Torque 6.1.2 and Maob 9.1.2. Bios update are done on most nodes. Configuration changes on GPFS done. Tech services CARNE code upgrade done.
|2018-03-16 1:00pm||2018-03-16 5:45pm||ISDA + NCSA OpenSource|
Security patches of VM servers as well as backend filesystem
Updates of Bamboo, JIRA, Confluence, BitBucket and CROWD
All systems will be unavailable for a brief period of time.
During updates of OpenSource services part of OpenSource will be offline for up to an hour.
|Updated fileserver (brief struggle with zfs and kernel updates). Updates of proxmox servers, Updated JIRA, Confluence, ButBucket and CROWD. Bamboo will be done later this weekend.|
Nebula Openstack cluster
|Security and filesystem patches||All instances and Nebula services were unavailable||Filesystem updates and security patches were applied. Filesystem is more responsive, but ~20 instances are repairing from problems that occurred before the outage.|
Lingering issues on select nodes following March PM
Following resolved by 13:23:
Resolved by 16:20:
|All systems were unavailable for maintenance.||Completed and most systems back online. Lingering issues for lsst-qserv-master01, lsst7, lsst-xfer, lsst-dts, and lsst-l1-cl-dmcs are being tracked in a separate status event.|
|Remote Access VPN||An issue with authentication for the VPN has occurred.||Any new connections will not be established. Existing connections are unaffected.||Authentication services were restored.|
|2018-03-09 10:08am||2018-03-09 11:00am||Campus Cluster||According to IBM, cc-mgmt1 was a culprit on halting communication across the cluster during the GPFS snapshot process.||User can't login or access to filesytem.||Rebooted cc-mgmt1 and restarted services (RM & Scheduler).|
|2018-03-09 06:05||2018-03-09 08:00||public-linux, www.ncsa.illinois.edu, & events.ncsa.illinois.edu||A routine kernel upgrade resulted in failure of the OpenAFS client on these servers.||OpenAFS storage was unavailable on these servers, resulting in the website failures.||Resolved. Packages were updated and OpenAFS reinstalled.|
|2018-03-07 15:00||2018-03-07 16:10||LSST||qserv-db12 had one failed drive in the OS mirror replaced but the other was presenting errors as well so the RAID could not rebuild. The Qserv system would have been unavailable during this maintenance.||qserv-db12||The node was taken down for replacement of the 2nd disk, to rebuild the RAID in the OS volume, and to reinstall the OS.|
|ESnet Peering||The connection servicing our direct peering with ESnet will be moved during this window.||Connections will be rerouted over a redundant peering. No service impact is expected.||The connection was successfully migrated and the peering with ESnet was brought back into service without issue.|
|WAN Connectivity Degraded||The router servicing several of our WAN connections is currently in a degraded state.||Traffic has been gracefully rerouted. No user facing connectivity issues have been reported.||Graceful failover to the backup routing engine cleared a fault condition and affected peerings were re-established.|
|2018-02-27 07:15||2018-02-27 09:10||Campus Cluster scheduler||Scheduler become unresponsive||Job submission & starting new jobs||Rebooted the node, restarted RM & Scheduler.|
|2018-02-26 06:00||2018-02-27 01:35||All Blue Waters Services||Security Patch CLE, SU26 Lustre patch||All Blue Waters resources are unavailable||Blue Waters returned to service at 1:35AM 27th Feb, with HPSS returned earlier at 10PM 26th Feb.|
|2018-02-23 16:30||2018-02-23 16:30||Kerberos Admin service||KDC configuration was modified to allow creation of service principles that can create and modify host and service principles.||kadmin service was unavailable for 1 second while new config was read.|
We can now delegate to group or users the ability to create and manage host keys and service principles.
|2018-02-23 08:00||2018-02-23 09:00||LSST Puppet Changes||Rolled out significant logic and organization of the Puppet resources in NCSA 3003 data center in order to standardize between LSST Puppet environments at NCSA. We had done extensive testing and did not expect any outages or disruption of services.|
No interruption of services.
Changes being applied to: lsst-dev01, lsst-dev-db, lsst-web, lsst-xfer, lsst-dts, lsst-demo, L1 test stand, DBB test stand, elastic test stand.
|Updated successfully with no interruption of availability or services.|
|2018-02-21 13:30||2018-02-22 00:39||ESnet 100G Peering Down||There is a suspected fiber cut between Urbana and Peoria on ICCN optical equipment. Our 100G direct WAN path to ESnet rides over this optical path and is thus currently down. The fiber vendor has identified the source of the problem (high water caused the fiber to be pulled out of a splice case)||Nothing. All traffic destined for ESnet or resources that would normally take the ESnet WAN path will reroute through our other WAN paths||Repaired.|
|2018-02-21 08:00||2018-02-21 20:00||Campus Cluster|
Campus Cluster February Maintenance
|All systems were unavailable|
Completed partially and following items are reschedule for next maintenance.
|2018-02-05 10:45||2018-02-21 13:30||ICCP Networking - Outbound|
A hardware failure on one of the two core switches for ICCP caused that switch to enter a degraded service mode and eventually fail completely. This was also combined with software bugs that caused looping of packets between the two cores in the MC-LAG. The other core was still functioning properly and was providing connectivity for all ICCP/ADS/DES systems normally for the duration of the degraded service time period. A hardware replacement RMA was initiated. The hardware came in but the hardware alone did not fix the issue. We then waited until a ICCP PM where we could test things without interruption of service and we upgraded the code and put in some bug mitigation configuration changes. These things combined solved the issues.
|Nothing as far as production. During the period where cc-core0 was down, aggregate bandwidth outbound was 40Gbps instead of the normal 80Ghps.||As of now the cores are both in production and stable.|
|2018-02-16 12:00||2018-02-16 12:30||IPSEC VPN||The appliance servicing various IPSEC VPN connections was patched.||Nothing||Patch was successful utilizing the failover capability of the VPN cluster to mitigate any service interruptions|
|2018-02-15 08:00||2018-02-15 13:00||LSST|
|All systems were unavailable.||Completed and all systems back online.|
|2018-02-13 08:00||2018-02-06 09:00||Certificate System Firewall 2||Upgrade software to current production version. No interruptions to service expected||CA services||FW upgraded - services were interrupted due to failed routing service.|
|2018-02-13 06:00||2018-02-13 06:30||AnyConnect VPN||Patches are being applied to the AnyConnect VPN appliance||Access to the NCSA AnyConnect VPN will be unavailable.||The VPN has been patched and client connections have been re-established.|
|2018-02-10 02:00||2018-02-10 10:35||Campus Cluster||GPFS snapshot hang and lock the filesystem||All systems were inaccessible. Lost running jobs.||Gather information for IBM, bounce the filesystem and reboot the cluster|
|2018-02-06 07:00||2018-02-06 17:35||iForge||Quarterly Maintenance (20180206 Maintenance for iForge)||All Systems were unavailable during the maintenance.||Planned maintenance completed successfully|
|2018-02-06 08:00||2018-02-06 09:00||Certificate System Firewall 1||Upgrade software to current production version. It is expected that current connections will be interrupted and a retry will be required.|
|2018-02-01 16:30||2018-02-01 16:45||sslvpn.ncsa.illinois.edu||We are rebooting our VPN appliances to mitigate a critical security vulnerability that allows for remote code execution exploits. That vulnerability is described here: https://tools.cisco.com/security/center/content/CiscoSecurityAdvisory/cisco-sa-20180129-asa1||Certain Industry partners' site-to-site VPNs||VPN rebooted without incident. Service was restored at 4:34PM.|
|2018-02-01 16:30||2018-02-01 16:45||vpn.ncsa.illinois.edu||We are rebooting our VPN appliances to mitigate a critical security vulnerability that allows for remote code execution exploits. That vulnerability is described here: https://tools.cisco.com/security/center/content/CiscoSecurityAdvisory/cisco-sa-20180129-asa1||Certain Industry partners' site-to-site VPNs and the NCSA remote access VPN service will be down during the maintenance. Any users connected to the NCSA VPN at the time of the maintenance will lose connectivity.||VPN rebooted without incident. Service was restored at 4:34PM.|
|2018-01-29 10:05||2018-01-29 10:10||LSST verify worker nodes and lsst-dev||A network flap on the LSST network caused GPFS ejection of some nodes. Network and security is investigating||a few of the LSST nodes for 2-5 minutes and 2 jobs||Qualys scan time frame changed and investigatino continues.|
|2018-01-29 12:27||2018-01-29 12:31||NCSA Jabber service||Jabber service was restarted to install a new SSL certificate.||NCSA Jabber was down momentarily||NCSA Jabber restarted with new SSL certificate|
|2018-01-26 13:00||2018-01-26 13:15||LSST NFS service slowdown||A cron for lenovo system cleanup was run, and caused the lenovo box to showdown services. The NFS service was starved.||lsst-dev NFS showed stale mounts||cron deleted, and re-written.|
|Wed 1/24/2018 13:35||Wed 1/24/2018 14:55||LSST NFS service||We were notified by NCSA security team that there was a stale NFS mount on one of the LSST test nodes. NFS services stopped working||All NFS mounts for LSST systems such as lsst-demo and lsst-SUI were not working||NFS server was rebooted.|
|Tue 1/23/2018 23:00||Wed 1/24/2018 01:25||Condo storage services||Hit a known bug in GPFS 188.8.131.52 for quota management.||All Condo services from 11pm to 1:25 am||Need to upgrade to a newer level of GPFS, but for now we have lowered frequency of the check_fileset_inodes script|
|2018-01-22 07:00||2018-01-22 13:05||Blue Waters Compute Nodes||Blue Waters compute nodes were bounced to resolve issues caused by previous home file system outage (due to bad OST)||Compute nodes were down, scheduler was paused.||Compute nodes were bounced successfully and returned to full service.|
|2018-01-21 08:42||2018-01-21 11:30||Netsec-vc switch stack - FPC 4||Switch member 4 of the netsec switch stack was down. Severe filesystem corruption occurred on the primary partition.||Any hosts connected to member 4 of that switch that were not redundantly connected to other switches in the stack.||The switch was repaired by doing a full reformat/reinstall of JunOS. Everything is back into production.|
|2018-01-20 22:00||2018-01-21 0300||Condo file systems||Bringing the Roger disk into the condo, commands executed from the Roger GPFS servers caused the cluster to arbitrate for GPFS servers.||All condo file systems mounted on nodes.||The SSH configuration was changed on the Roger GPFS servers to include the Condo GPFS server IP's. All file systems were returned to normal with no other problems and no remounts required.|
|2018-01-18 17:00||2018-01-19 15:00||ISDA Hypervisors, NCSA Open Source||Hypervisor updates.||All systems were down for short amount of times as hypervisors rebooted||All patches applied.|
|2018-01-18 00:00||2018-01-18 24:00||Campus Cluster||Copying all data to new filesystem. Deploying new Storage (14K). Dividing cluster into two (IB & Ethernet). Upgrading GPFS to 184.108.40.206. Deploying new management node and new image server (if time permit). Applying Security patches to compute nodes(no FW update at this time).||All systems unavailable.||New Storage System was brought online, additional capacity and performance was added.|
|LSST||LSST Firewall outage in NPCF. Both pfSense firewalls were accidentally powered off.|
PDAC (Qserv & SUI) and verification clusters were inaccessible, as well as introducing GPFS issues across many services, e.g. lsst-dev01.
|The pfSense firewall appliances were power cycled and services restored.|
|2018-01-18 12:58||2018-01-18 14:10||Code42 Crashplan backup system||Code42 Crashplan server were upgraded to latest JDK and Code42 6.5.2.||Clients were unable to perform restores or push files into backup archive from roughly 13:35 - 13:55||Code42 servers are now running latest security updates to the crashplan service.|
|2018-01-18 08:00||2018-01-18 10:00||LSST||Monthly OS updates, network switch updates, firmware updates, etc.||All dev systems unavailable. Qserv and SUI nodes will remain available.|
|2018-01-17 10:35||2018-01-17 13:00||RSA Authentication Manager Servers||Upgraded to Authentication Manager 8.1sp1p7||No systems should have seen any impact||Latest security patches are applied.|
|2018-01-12 06:00||2018-01-12 10:00||Decommission NCSA Rocket.chat||The old NCSA Rocket.chat service was shutdown.|
Any archived conversations or content are no longer be available to users.
|NCSA Rocket.chat service was shutdown and redirected to NCSA @ Illinois Slack.|
|Friday, Jan 12th, 0000-0600 CST||Internet2||Engineers from Internet2 will be migrating our BGP peering with I2's Commercial Peering Service (CPS) to a new location. Small disruptions may occur with the maintenance for the CPS service, but no user traffic disruptions should occur.||None, Alternatives routes are present.||none||Maintenance was completed successfully.|
|2018-01-11 08:00||2018-01-11 13:30||LSST||Critical patches on lsst-dev systems (incl. kernel updates).||All systems unavailable.|
|Thursday, Jan 11th, 0000 CST||Thursday, Jan 11th, 0400 CST|
Connectivity to Internet2 and backup LHCONE peerings - ICCP and MWT2 respectively
|Engineers from Internet2 performed maintenance that affected certain BGP peerings that exist on the device that is ICCP/MWT2's upstream router, CARNE. Specifically, both the 100G Internet2 peering and the Internet2 LHCONE peering on CARNE were disrupted during this timeframe. MWT2 currently gets to LHCONE through CARNE's ESnet peering, which was fully functional. They also were able to get to UChicago through CARNE's OmniPoP 100G peering. As for ICCP, traffic to/from Internet2 based routes rerouted through the ICCN.||Nothing was reported to be service impacting by this maintenance from neither ICCP nor MWT2.||Successful maintenance was completed.|
|2018-01-08 10:47||2018-01-08 11:30||Nebula||Storage nodes lost networking||All nebula instances||Storage nodes were brought back online, instances were rebooted|
|2018-01-05 17:00||Nebula||Nebula was shut down for hardware and software maintenance from January 2nd, 2018 at 9am until January 5th, 2018 at 5pm. Spectre and Meltdown patches were applied, as well as all firmware updates, OS/distribution updates, and the filesystem was upgraded.||All systems were unavailable.||Faster system that is now homogenous, so OpenStack upgrades are now possible.|
|2018-01-04 17:00||2018-01-05 20:00||Blue Waters||One OST hosting the home file system has three drives failed simultaneously.||Portion of home file system (with data on the affected OST) are not accessible.|
Repair works were carried out on the failed OST. Scheduler continued to operate but restricting only jobs not affected by the failed OST to start.
Full operation resumed after successful recovery of the failed OST.
|2017-12-20 08:00||2017-12-20 10:00||LSST||(1) Firewall maintenance (08:00-09:00) and (2) migration of NFS services (08:00-10:00).|
Firewall maintenance: There should be no noticeable effect but scope of service includes most systems at NPCF (including PDAC, SUI, and Slurm/batch/verify nodes).
Migration of NFS services: SUI and lsst-demo* nodes.
|Maintenance completed without issues.|
|2017-12-14 06:00||2017-12-14 20:30||LSST||Monthly OS updates, network switch updates, firmware updates, etc.||All systems unavailable.|
All systems back online.
|2017-12-13 09:00||2017-12-13 11:00||JIRA Upgrade||Upgraded JIRA to version 7.6 from 7.0||NCSA Jira||Succesfully upgraded|
|2017-12-13 06:30||2017-12-13 07:39||NCSA Jabber||Attempted to upgrade Openfire XMMP jabber software.||NCSA Jabber was unavailable during the upgrade.||The upgrade failed. Jabber is available, but still running the old version. The upgrade will be rescheduled.|
|2017-12-11 10:00||2017-12-11 16:00||Unused AFS fileserver were upgraded to 1.6.22||After moving all volumes to servers updated on 2017-12-07, the now unused AFS servers were upgraded to OpenAFS 1.6.22.||No impact to other systems as they were unused at the time they were upgraded.||All of NCSA's afs cell is running on OpenAFS 1.6.22|
|2017-12-09 03:00||2017-12-09 07:42||BlueWaters Portal||The BlueWaters portal software crashed. Automated monitoring processes did not restart it correctly.||The BlueWaters portal website was unavailable.||The BlueWaters portal service was manually restarted and the website is available.|
|2017-12-09 1000hrs||2017-12-09 1400hrs||Globus Online (Globus.org)||Please be advised that the Globus service will be unavailable on Saturday, December 9, 2017, between 10:00am and 2:00pm CST while we conduct scheduled upgrades. Active file transfers will be suspended during this time and they will resume when the Globus service is restored. Users trying to access the service at globus.org (or on your institution's branded Globus website) will see a maintenance page until the service is restored.||All NCSA Globus endpoints.|
|2017-12-07||2017-12-07||Unused AFS file servers were upgraded to 1.6.22||Three unused AFS fileserver were upgraded to the latest 1.6.22 release of OpenAFS||No impact to other systems as they were unused.||These AFS fileserver can no longer be crashed by malicious clients.|
|2017-12-07||2017-12-07||AFS database servers were upgraded to 1.6.22||The three database servers were upgraded to the latest 1.6.22 release of OpenAFS||No modern clients noticed the staggered updates.||These servers can no longer be crashed by malicious clients.|
|2017-12-05 16.00||2017-12-05 16:20||dhcp.ncsa.illinois.edu||NCSA Neteng will be migrating the DHCP server VM to Security team's VMware infrastructure.|
- Hosts on the NCSAnet wireless network might be impacted.
|Maintenance was completed successfully and services are running as expected.|
|2017-12-02 09:30||2017-12-02 11:45||NCSA opensource||Upgrade of Bamboo, JIRA, Confluence, BitBucket FishEye, and CROWD||Sub services of opensource can be down for a short time.||All services upgraded and running as normal.|
|ROGER OpenStack cluster||I/O issues highlighted that GPFS CES NFS servers probably shouldn't run 400+ days without reboot||ROGER's OpenStack and the various services which were hosted therein, including JupyterHub Server||reboot of all nodes, including CES servers as well as the reboot of all hypervisors (with the fallout being one node required fsck and second reboot and another node/hypervisor is still unavailable) cleared most of the problems. I/O contention was felt as many instances were simultaneously attempting to start/restart. instances that were housed on the unavailable node are being migrated to another hypervisor|
Update the fileserver that hosts VM's
NCSA Open Source unavailable
Network issues delayed updates
|2017-11-21 16:00||2017-11-21 16:40||Code42 Crashplan||The Code42 crashplan infrastructure was upgraded to version 6.5.1 to apply security and performance improvements||Clients transparently reconnected to servers after they restarted||Now running on Code42 version 6.5.1|
|2017-11-20 9:00||2017-11-20 16:38||Nebula Openstack cluster||Nebula OpenStack cluster was unavailable for emergency hardware maintenance. A failing RAID controller from one of the storage nodes and a network switch were replaced.||Not all instances were impacted. Running Nebula instances that were affected by the outage were shut down, then restarted again after we finished maintenance.|
Nebula is available.
|2017-11-16 16:46||2017-11-20 12:40||NCSA JIRA||JIRA wasn't importing some email requests properly after the NCSA MySQL restart.||Some email sent to JIRA via help+ addresses wasn't being imported.||JIRA is now accepting email and all email sent while it was broken has now been imported as expected.|
BW LDAP Master
|Scheduled maintenance||Updated LDAP lustre quotas to bytes and add archive quotas. IDDS will track and drive quota changes with acctd.||Production continued w/o interruption. BW LDAP master was isolated, lustre quotas changed to bytes with the addition of archive quotas. Replicas pulled updates w/o error.|
|2017-11-16 14:30||2017-11-16 16:52||Internal website (MIS Savanah)||A database table used by MIS tools became corrupted.||The website would become unresponsive every time the corrupted database table was accessed.||OS kernel and packages where updated during debugging. The MIS database table was restored and the website came back online.|
|2017-11-16 16:46||2017-11-16 16:48||NCSA MySQL||The NCSA MySQL server had to be restarted in order to delete the corrupted table used by MIS.||All services that use MySQL were down during the outage. This includes: Confluence, JIRA, RT, and lots of websites||MySQL was restarted successfully.|
|2017-11-16 0800||2017-11-16 1200||LSST||Monthly OS updates, plus first round of Puppet technical debt changes (upgrading to best design & coding practices)|
All systems unavailable from 0800 - 1000 hrs.
GPFS unavailable from 0800 - 1000 hrs.
PDAC systems unavailable from 0800 - 1200 hrs.
Completed. OS kernel and package updates. Slurm upgrade to 17.02.
|2017-11-15 13:30||2017-11-15 15:10||RSA Authentication Manager||RSA Authentication Manager were patched to fix cross site scripting vulnerabilities and other fixes||Nothing was affected by the update||RSA Authentication Manager is running 8.2 SP1 P6. Process worked as expected.|
|2017-11-15 - 13:30||2017-11-15 - 14:30||BW 10.5 Firewall Upgrade Part 2||The normal active, "A" unit, NCSA BW 10.5 Firewall will be upgraded and then normal fail-over status will be re-enabled.||The possibility of connection resets when the A unit comes back from being upgraded and state is being sycned.||Completed, process worked as expected.|
|2017-11-14 11:27||2017-11-14 11:33||LDAP||LDAP was unresponsive to requests.||Several services hung while authentication was unavailable.||LDAP services were killed and restarted.|
|ROGER Hadoop/Ambari||cg-hm12 and cg-hm13 took minor disk failures which crashed the node||Ambari was effectively off-line||rebooted node, and node ran fsck as part of its startup sequence, node booted properly|
|ROGER hadoop/ambari||hard drive failures on cg-hm10 and cg-hm17||certain ambari services and HDFS||cg-hm17 returned to service after power cycle and reboot, cg-hm10's hard drive didn't respond to a reboot|
|2017-11-11 16:58||2017-11-11 19:09||Blue Waters||Water leak from XDP4-8 causing high temperature to c12-7 and c14-7.||EPO on c12-7 and c14-7.||Scheduler was paused to place system reservations on compute nodes in affected cabinets, then resumed.|
|2017-11-10 14:00||2017-11-10 14:45||NCSA Open Source||Upgrade of the following software: Bamboo, JIRA, Confluence, and BitBucket||Updates will happen in place and will result in minimal downtime of components.||completed, minimal interruption of service|
|2017-11-10 - 08:00||2017-11-10 - 08:30||CA Firewall Upgrade - B unit||the stand-by, "B" unit, NCSA Certificate Service Firewall will be upgraded to same version as A unit.||Expect no impact to services||completed, no interruption of service|
|Netdot||netdot.ncsa.illinois.edu was migrated to Security's VMware infrastructure.||During the downtime users weren't able to activate or deactivate their network connections via Netact.||Migrated successfully. Netdot is up and running.|
|2017-11-08 06:00||2017-11-08 15:00||ITS vSphere vCenter||ITS vSphere was upgraded to the latest version of VMware vCenter. New access restrictions were also be put into place.||All VMs remained online during the maintenance, but management through vCenter was offline during the upgrade.||Upgrade completed successfully.|
|2017-11-08 09:30||2017-11-08 10:00||BW 10.5 Firewall Upgrade Part 1||the stand-by, "B" unit, NCSA BW 10.5 Firewall will be upgraded and then traffic redirected through it for load testing before the "A" unit is upgraded||Expect no impact to services||Upgrade completed successfully. Some states were reset when traffic switched to the B unit.|
|2017-11-07 7:00||2017-11-07 18:37||iForge|
Update OS image.
|iForge (and associated clusters)|
All production systems are back in service
|2017-11-07 - 13:30||2017-11-07 - 15:00||CA Firewall Upgrade Part 2||The normal active, "A" unit, NCSA Certificate Service Firewall will be upgraded and then normal fail-over status will be re-enabled.||The possibility of connection resets when the A unit comes back from being upgraded and state is being sycned.||Completed upgrade|
|2017-11-06 15:28||2017-11-06 15:53||Blue Waters||EPO happened to c12-7 and c14-7.||HSN quiesced.||Scheduler was paused to place system reservations on compute nodes in affected cabinets, then resumed.|
|2017-11-03 16:21||2017-11-03 16:32||LDAP||LDAP was unresponsive to requests.||Several services hung while authentication was unavailable.||LDAP services was killed and restarted.|
|2017-11-02 09:00||2017-11-02 16:00||LSST||LSST had a GPFS server that was down and had failed over to the other server for NFS.||The GPFS client’s failed over automatically, and we manually failed over the NFS in the morning.||NFS exports were moved to an independent server. IBM was at NCSA and is continuing to debug the problems.|
|2017-10-31 17:11||2017-11-01 11:13||LSST||GPFS degraded/outage|
most NCSA-hosted LSST resources experienced degraded GPFS performance
hosts with native mounts (PDAC) experienced an outage
|A deadlock at 17:11 yesterday temporarily caused slow performance. Then one GPFS server went offline at 18:21 and services failed over. NFS mounts (qserv/sui) were reported as hanging by a user at 09:12 today but may have been degraded over night. Affected nodes were rebooted and NFS mounts recovered by 11:13. IBM is onsite diagnosing issues with the GPFS system and ordering repairs (including a network card on one server).|
|2017-10-31 15:30||2017-10-31 16:00||LSST||GPFS outage|
most NCSA-hosted LSST resources
native mounts (e.g., lsst-dev01, verify-worker*) and NFS mounts (e.g., PDAC)
|All disks in the GPFS storage system went offline temporarily and came back online by themselves. NFS services were restarted. Client nodes all recovered their mounts on their own. Logs have been sent to the vendor for analysis.|
|2017-10-31 - 13:30||2017-10-31 - 14:30||CA Firewall Upgrade Part 1||the stand-by, "B" unit, NCSA Certificate Service Firewall will be upgraded and then traffic redirected through it for load testing before the "A" unit is upgraded||Expect no impact to services||Upgrade completed successfully. Some states were reset when traffic switched to the B unit.|
|2017-10-30 18:36||2017-10-31 00:46||LSST||GPFS outage|
most NCSA-hosted LSST resources
native mounts (e.g., lsst-dev01, verify-worker*) and NFS mounts (e.g., PDAC)
|GPFS servers were rebooted. lsst-dev01 and most of the qserv-db nodes were also rebooted. Native GPFS and NFS mounts were recovered. May have been (unintentionally) caused by user processes but will continue to investigate..|
|2017-10-25 22:00||2017-10-26 11:20||LSST||full/partial GPFS outage|
full outage for GPFS during 22:00 hour on 2017-10-25
outage for NFS sharing of GPFS (for qserv, sui) continued through the night
full outage for GPFS recurred 2017-10-26 around 08:44
|All GPFS services and mounts have been restored.|
|2017-10-26 09:04||2017-10-26 09:04||Various buildings across campus, including NPCF and NCSA||Issue with an Ameren line from Mahomet caused a bump/drop/surge in power that lasted 2ms||LSST had approximately 20 servers at both NPCF and NCSA buildings reboot||Was a momentary issue with minimal effect to most systems|
|2017-10-26 00:00||2017-10-26 08:00||ICCP||gpfs_scratch01 was filled by a very active user||Additional space in scratch wasn't available||Out of cadence purge was run to free 2TB, users jobs held in scheduler; user contacted|
|2017-10-25 06:00||2017-10-25 14:05||Blue Waters||Security Patching of CVE-2017-1000253 security vulnerability.||Restricted access to logins, scheduler and compute nodes. HPSS and IE nodes are not affected.||System was patched. Logins hosts are made available at 9am. The full system is returned to service at 14:05.|
|2017-10-24 09:50||2017-10-24 20:10||LSST||Network outage / GPFS outage|
All LSST nodes from NCSA 3003 (e.g., lsst-dev01/lsst-dev7) and NCPF (verify-worker, PDAC) that connect to GPFS (as GPFS or NFS) lost their connections.
All LSST nodes at NPCF lost network during network stack troubleshooting and replacement of 3rd bad switch.
|A 3rd bad switch was discovered and replaced. All nodes have network and GPFS connectivity once again.|
|2017-10-23 08:00||2017-10-24 05:00||Campus Cluster||Campus Cluster October maintenance.||Total outage of the cluster.||Replaced core ethernet switches from share services pod. Run new ethernet cables for share services pod. Moved DES rack from share services pod to ethernet only pod. Deployed new image with patched.|
|LSST||First one then two public/protected network switches went down in racks N76, O76 at NPCF||Mostly qserv-db[11-20] and verify-worker[25-48]; there was also shorter outage for qserv-master01, qserv-dax01, qserv-db[01-10], all of SUI, and the rest of the verify-worker nodes.||Two temporary replacement switches were swapped in. Maintenance and/or longer-term replacement switches is being procured for the original switches.|
|2017-10-18 13:00||2017-10-18 14:00||Networking||Replaced a linecard in one of our core switches due to hardware failure.||Any downstream switches were routed through the other core switch.||All work was completed successfully.|
|2017-10-19 08:00||2017-10-19 21:30||LSST||Outage and migration of qserv-master01: provisioning of new hardware, copying of data from old server to new.||qserv-master01 (and any services that depend on qserv-master01, which may include services provided by qserv-db*, qserv-dax01, and sui*)|
UPDATE (2017-10-19 15:15) OS install took much longer than anticipated, completed at 15:00. Data sync is started. Extending outage till 22:00.
|10-19 08:00||2017-10-19 12:00||LSST||Routine patching and reboots, pfSense firmware updates (NPCF), Dell server firmware updates (NPCF).||All NCSA-hosted resources except for Nebula.||Maintenance completed successfully. (qserv-master migration is ongoing, see separate status entry)|
|2017-10-18 14:45||2017-10-18 15:35||Campus Cluster||Restart of resource manager failed after removing all block array jobs.||Job submission||Opened case with Adaptive (#25796). Found more array jobs and bad jobs in jobs directories. Removed all of those.|
|2017-10-15 08:15||2017-10-15 08:30||Open Source||Emergency upgrade of Atlassian Bamboo.||Bamboo will be down for a few minutes during this outage window.||Bamboo upgraded to the latest version.|
|2017-10-14 22:15||2017-10-14 23:35||Campus Cluster||Scheduler crash||Job submission||Opened case with Adaptive, run diag and uploaded the output along with the core file. Restarted the moab.|
|2017-10-14 13:00||2017-10-14 15:23||Campus Cluster||Resource manager crash||Job submission||Applied patch from Adaptive, which help with faster recovery. Suspend/block all current and new array jobs until we have a resolution.|
|2017-10-06 09:00||2017-10-11 01:00||Nebula||Gluster and network issues|
1) Gluster sync issues continue from 2017-10-05's Nebula incident.
All instances have been restarted and are in a state for admins to run. Some mounted file systems might require a fsk to verify. If there are other issues please send a ticket.
As the file system continues to heal we may see slower interaction.
|2017-10-10 16:30||2017-10-10 19:10||Campus Cluster||Resource manager crash||Job submission||After removing problematic jobs from queue and we were able to restart the RM. Opened the case with Adaptive and forwarded those job scripts and core files.|
|2017-10-05 14:00||2017-10-05 17:00||Nebula||Gluster sync issues||One of the gluster storage servers within Nebula had to be restarted.||Approximately 100 VM instances experienced IO issues and were restarted.|
|2017-10-06 08:00||2017-10-06 17:00||NCSA direct peering with ESnet|
A fiber cut between Peoria and Bloomington caused our ESnet direct peering to go down.
|All traffic that would have taken the ESnet peering rerouted through our other WAN peers. As such there were no reported outages of connectivity to resources that users would normally access via this peering||The fiber cut has been repaired and the peering has been re-established.|
|2017-10-06 08:00||2017-10-06 10:00||LSST||Kernel and package updates to address various security vulnerabilities, including the PIE kernel vulnerability described in CVE-2017-1000253. This will involve an upgrade to CentOS 7.4 and updates to GPFS client software on relevant nodes.||All NCSA-hosted LSST resources except for Nebula (incl. LSST-Dev, PDAC, and verification/batch nodes) will be patched and rebooted.||Maintenance completed successfully. Pending updates to a couple of management nodes (adm01 and repos01) and one Slurm node that is draining (verify-worker11).|
|2017-10-4 07:40||2017-10-4 09:55||Campus Cluster||Resource Manager crash||Job submission||Failure on initial restart attempt. After looking through the core, decided to try a restart again without any change. This time it worked.|
|2017-10-03 13:00||2017-10-03 19:00||Campus Cluster||Resource Manager crash||Job submission||After removing ~30 problematic jobs from queue and we were able to restart the RM. Opened the case with Adaptive and forwarded those job scripts and core files.|
|2017-09-21 02:57||2017-09-21 09:40||Storage server (AFS, iSCSI, web, etc)|
The parchment storage server stopped responding on the network.
|Replaced optical transceiver on the machine and networking restarted. Also updated kernel and AFS.|
|2017-09-20 08:00||2017-09-20 13:45||Campus Cluster||September Maintenance||Total cluster outage||Maintenance completed successfully.|
|2017-09-20 08:00||2017-09-20 11:30||NCSA Storage Condo ||Normal maintenance --Firmware upgrade on Netapps so new disk trays could be attached for DSIL||total file system outage||The quarterly maintenance was complete|
|2017-09-18 11:20||2017-09-18 13:30||Active Data Storage||RAID Failure in NSD server and disk failure on secondary NSD server.||ADS service was unavailable||Recovered RAID configuration on NSD server and replaced failed disk on secondary NSD. ADS restored.|
|2017-09-15 06:20||2017-09-15 09:28||public-linux||OpenAFS storage was not running or mounted after rebooting to a new kernel.||AFS storage was not available from this server||Reinstalled the dkms-openafs package restarted the openafs-client. AFS is now working as expected.|
|2017-09-10 09:45||2017-09-10 11:30||NCSA Open Source||Upgrade of Bamboo, JIRA, Confluence, BitBucket, FishEye, Crowd||During the upgrade the services will be unavailable for a short amount of time.||All services upgraded successfully.|
|2017-08-31 11:07||2017-08-31 11:11||NCSA LDAP||NCSA LDAP Timeouts||NCSA LDAP was overloaded and timing out. Users were not able to authenticate via NCSA LDAP during that time.||NCSA LDAP stopped timing out at 11:11 am and authentication resumed.|
|2017-08-28 11:55||2017-08-28 12:59||NCSA GitLab||NCSA GitLab server ran out of disk space for the OS||The web interface at https://git.ncsa.illinois.edu wasn't working||Web interface is now working. Space freed up by clearing CrashPlan caches.|
|2017-08-24 13:00||2017-08-24 14:30||netact.ncsa.illinois.edu||Transient config issues from some system patching caused apache to not be able to start on the netact server||Network Activation||The issues were fixed and Network Activation is working again|
|2017-08-24 08:00||2017-08-24 15:30||LSST||Rack upgrades in NCSA 3003||Most LSST Developer services offline during upgrade||All LSST systems are back online with new racks and switches|
|2017-08-24 08:00||2017-08-24 09:30||LSST||monthly maintenance for NPCF (includes patching to address CESA-2017:1789 and CESA-2017:1793)||adm01, backup01, bastion01, monitor01, object*, qserv*, sui*, verify-worker*, test0*||Maintenance was successfully completed.|
|2017-08-23 09:21||2017-08-23 16;50||aForge/iForge||gpfs failed during an upgrade of GPFS on the iforge storage nodes. There was an IB hiccup at the time, but causality is unclear||all jobs on iforge were aborted, gpfs clients needed to be upgrade, all gpfs client nodes were rebooted||iForge went production shortly before 5:12pm. aforge went "production" at ~1630|
|2017-08-22 20:00||2017-08-22 30:00||Patching DHCP service||Patching OS and services on DHCP1.||Will need to reboot DHCP server a few times during this process. During the time dhcp will be unavailable. This is during the evening so I don't expect any direct issues from this.||Patching has been completed.|
|2017-08-16 08:00||2017-08-16 16:00||Campus Cluster||August Maintenance||Scheduler and resource manager down||Upgraded Moab 9.1.1 and Torque 6.1.1.|
|2017-08-16 08:00||2017-08-16 09:15||None||Replace Line Card in Core Switch||I believe all systems connected to this switch, are multihomed and will not experience an outage.||The line was has been successfully replaced.|
|2017-08-16 00:30||2017-08-16 02:30||Blue Waters||Two cabinets (c10 & c11) had EPO due to XDP control valve failure.||Scheduler was paused to isolate failing parts, resumed at 2:09.||Parts replaced and cabinets were returned to service.|
|2017-08-08 7:00||2017-08-09 3:00||iforge/cfdforge/aforge|
Update OS image to RH 6.9
Update GPFS to version 4.2.3-2
Redistribute power drops
|All four clusters were updated.|
All items on checklist completed.
|2017-08-03 06:45||2017-08-03 07:35||NCSA Jabber upgrade||Upgraded Openfire XMMP jabber software||NCSA Jabber was unavailable during the upgrade.||Jabber was upgraded to the latest version of Openfire|
|2017-07-28 17:00||2017-07-31 evening||Update - All of the production data has been migrated except for the largest object table. That is loading now, then the user space will be loaded. Should all hopefully be done by this evening. Migration of operational database to new hardware happening during the weekend.||DES old operational database||migration done successfully. Some other maintenance tasks that will give DES additional disk space was done, too and some performance improvements.|
|2017-07-27 11:00||2017-07-28 15:00||netact.ncsa.illinois.edu|
The netact.ncsa.illinois.edu network activation server VM needed to be restored from backup
|Network Activation service||The service has been fully restored|
|2017-07-25 02:36||2017-07-25 18:00||Campus Cluster / Scheduler down||Blip on mgmt1 causing GPFS drop and scheduler to crash||Scheduler offline||Still taking long time for Scheduler to initialize but jobs can start and run as usual. Opened case with Adaptive.|
|2017-07-20 09:00||2017-07-20 17:00||ROGER Ambari and OpenStack||Updates to openstack control node and the Ambari cluster||Ambari nodes (cg-hm08 - cg-hm18), OpenStack instances and servers||Openstack was back in service on time. Ambari had issues mounting hdfs was held out of service. HDFS was remounted on 25 July|
|2017-07-20 06:00||2017-07-20 10:00||All NCSA hosted LSST resources||Monthly OS patches (addressing issues including CESA-2017:1615 and CESA-2017:1680). Roll-out updated puppet modules. Batch nodes updated firmware.||All nodes in NCSA 3003 and NPCF (batch nodes) will reboot.||Overall success. Exceptions: verify-worker31 failed a firmware update and is out of comission (LSST-914) and there are connectivity issues for some VMs used by the NCSA DM team (IHS-365). adm01, backup01, and test[09-10] will be patched in the near future.|
|2017-07-19 08:00||2017-07-19 14:44||Campus Cluster||July Maintenance (applied security patch)||Cluster wide, except mwt2 nodes||Applied new kernel, glibc, bind patches and newest NVIDIA driver.|
|2017-06-30 0000||Blue Waters||Emergency maintenance to apply security patch addressing Stack Guard security vulnerability.||Compute, Login, Scheduler are offline.||Kernel and glibc library patched on all affected system.|
|2017-06-22 0800||2017-06-22 1200||All NCSA hosted LSST resources||CRITICAL kernel and package updates to address Stack Guard Page security vulnerability.|
Systems will be patched and rebooted.
|Outage was extended to last past 1000 until 1200. Systems were successfully patched as planned except for qserv-db12 and qserv-db27, which will not boot. We will follow up on those with a ticket.|
|2017-06-22 0800||2017-06-22 0930||LSST cluster nodes (verify-worker*, qserv*, sui*, bastion01, test*, backup01)||Deploy Unbound (local caching DNS resolver)||DNS resolving may have a short (~30 mins) delay.||Successfully deployed and all tests (including reverse DNS and intra-cluster SSH) pass.|
|Bluewaters||XDP shutting down causing EPO on cabinet c1-7 and c2-7.||Scheduler was paused to isolate the failing components, then resumed.||Warmswap of failing components, and returned them to service.|
|NCSA Open Source|
Security upgrade needed for Bamboo, will also update the following components: Bamboo, JIRA, Confluence, BitBucket, FishEye
|Most of the subcomponents of NCSA opensource will be down for a short time when the software is updated.||Upgraded Bamboo, JIRA, Confluence, BitBucket, FishEye to latest versions|
|ROGER Openstack nfs backend failed and was restarted||The primary CES server for the openstack backend failed and tried to fail over to the secondary server, which also failed. SET was notified and they had the CES nfs service back up by 1100||The RoGER openstack dashboard went down and needed a restart. Several VM's experienced "virtual drive errors" and will need to be restarted||SET is still investigating the cause of the GPFS CES service failover. CyberGIS is working with their users to get the affected VM's restarted|
|2017-06-15 0800||2017-06-15 0930||LSST cluster nodes (verify-worker*, qserv*, sui*, bastion01, test*, backup01)||Deploy unbound||DNS resolving may have a short (~30 mins) delay.|
Updates deployed successfully via new puppet module. All tests passed.
EDIT 2017-06-15 1500 - Reverse DNS not working, which broke ssh to qserv* nodes. Disaabled unbound.
|Network Core Switch||Network Engineering will be replacing a line card in one of our Core switches due to hardware issue.||All services should remain active. Any affected switch will have a second redundant link to the other core to pass traffic.||Line card was successfully replaced.|
|2017-06-08 12:00||2017-06-11 22:20||Campus Cluster (scheduler paused)||Disk Enclosure 3 failure on DDN 10K.||Lost redundancy and force us to drain the cluster.||Repair/replacement for controller can be time consuming so we took action to rebalance data out of failed enclosure. Scheduler was resumed as of 22:00.|
|2017-06-07 12:42||NCSA LDAP||The NCSA LDAP service crashed||NCSA LDAP service was unavailable||LDAP software and OS were updated and server rebooted. LDAP is working normally.|
|2017-05-31 20:06||2017-05-31 20:36||NCSA LDAP||The NCSA LDAP service was timing out||NCSA LDAP service was unavailable||The root cause of LDAP timeouts is still being investigated.|
|2017-05-22||2017-05-26||Campus Cluster VMs||Network issue ESXI (hypervisor) Boxes after maintenance||Could no longer able to login to start VMs. License Server, nagios, all MWT2 VMs were down|
The issue is fixed on 5/24. Restored license and Nagios service on 5/24. Moved MWT2 VMs to Campus Farm. All VMs return to service as of noon 5/26.
|5/12/2017||5/18/2017||Condo/NFS partitions only||the NFS partition for the condo became extremely unstable after a replication (normal daily maintenance) was completed. Many iterations with FSCK and IBM on the phone got it resolved, and then 1.5 days restoring files that had been put in Lost and found.||UofI library was switched to the READONLY version on the ADS during this time||The root cause is still being investigated.|
|2017-05-23 14:05||2017-05-23 14:13||NCSA LDAP||The NCSA LDAP service was timing out||NCSA LDAP service was unavailable||The issue is still being investigated, but seems to be steadily available since the incident.|
|2017-05-22 15:41||2017-05-22 15:51||idp.ncsa.illinois.edu|
|Apache Tomcat out of memory||InCommon/SAML IdP and OIDC authentication services were unavailable.||Service restored by failing over to secondary server while memory is being increased on primary server.|
|DES nodes on Campus Cluster||Could not communicate outside the switch||All nodes connected to switch in POD22 Rack2 @ACB||Upgraded the code on the switch resolved the issue.|
|05/20/2017 05:00||05/20/2017 21:09||Campus Cluster and Active Data Storage (ADS)||Total power outage at ACB||All systems currently reside at ACB|
Power was restored around 13:00hrs. We rotated ADS rack to align with Campus Cluster Storage Rack. Changed couple of VLAN IDs to reflect campus for future merger. ESXI boxes are down due to a configuration error after reboot. No major issue from output of FSCK from scratch02.
|05/17/2017 02:00||05/17/2017 10:45||Internet2 WAN connectivity||Intermittent WAN connectivity. The outage was a result of Tech Services' DWDM system, which provides us with our physical optical path up to Chicago via the ICCN. Specifically, the Adva card that our 100G wave is on was seeing strange errors, which was causing input framing errors for traffic coming in on this interface.||General WAN connectivity to XSEDE sites, certain commodity routes, and other I2 AL2S connections.||The Adva card was rebooted and we stopped seeing the input framing errors. Tech Services is working with Adva to find the root cause of the issues on the card.|
|5/11/2017||5/12/2017||ESnet 100G connection||NCSA and ESnet will be moving their 100G connection to a different location in Chicago.||We have several diverse high speed paths to ESnet and DOE, traffic will be redirected to a secondary path.|
|NCSA Jabber upgrade||Upgraded Openfire XMMP jabber software||NCSA Jabber was unavailable during the upgrade.||Jabber was upgraded to the latest version of Openfire|
|iForge, GPFS, License Servers||iForge Planned Maintenance||iForge systems, including the ability to submit/run jobs.||Pm was completed early at 1815|
|2017-05-06 22:00||2017-05-06 23:00||NCSA Open Source||Upgrades of Atlassian software||NCSA Open Source BitBucket||BitBucket is upgraded.|
|2017-05-06 09:00||2017-05-06 10:00||NCSA Open Source||Upgrade of Atlassian Software||Most services hosted at NCSA Open Source were down for 5 minutes during rolling upgrades.||The following services were upgraded: HipChat, Bamboo, JIRA, Confluence, FishEye and CROWD.|
|2017-05-05 17:43||2017-05-05 20:02||ITS vSphere||A VM node panicked||Several VMs died when the node panicked and were restarted on other VM nodes. This included LDAP, JIRA, Help/RT, SMTP, Identity, and others.||All affected VMs were restarted on other VM nodes. Most restarted automatically.|
|2017-04-27 18:10||2017-04-27 18:55||Campus Cluster||Another GPFS interruption||Both Resource Manager and Scheduler went down along with hand full of compute nodes.||Restarted the RM and Scheduler and rebooted all down nodes.|
|2017-04-27 13:11||2017-04-27 14:20||Nebula||glusterfs crashed due to this bug, so no instances could access their filesystems||All instances running on Nebula||Needed to reboot the node that systems were mounting from, but took the opportunity to upgrade all gluster clients on other systems while waiting for a reboot. Version 3.10.1 fixes the bug. All instances with errors in their logs were restarted.|
|2017-04-27 11:20||2017-04-27 12:45||Campus Cluster||GPFS interruption||Both Resource Manager and Scheduler went down.||Torque serverdb file was corrupted. Restore the file from this morning snapshot and modified the data to match the current state.|
|2017-04-26 12:00||2017-04-26 18:30||Condo||A bug in the delete of a disk partition from GPFS. a problem within GPFS||DES, Condo partitions, and UofI Library.||Partitions had been up for 274 days, and many changes. The delete partition bug caused us to stop ALL operations on the condo and repair each disk through GPFS. Must have quarterly maintenance. Just too complicated to go a year without reseting things.|
|2017-04-19 16:54||2017-04-20 08:45||gpfs01, iforge|
Filled-up metadata disks on I\O servers caused failures on gpfs01.
|iforge clusters, including all currently running jobs.|
Scheduling on iForge was paused for the duration of the incident. Running jobs were killed.13% metadata space was freed. Clusters were rebooted and scheduling resumed.
|2017-04-19 08:00||2017-04-19 13:00||Campus Cluster||Merging xpacc data and /usr/local back to data01 (April PM)||Resource manager and Scheduler were unavailable during the maintenance.||Once again, /usr/local, /projects/xpacc and /home/<xpacc users> are mounting from data01. No more split cluster.|
|2017-04-04 (1330)||2017-04-04 (1600)||Networking||Some fiber cuts caused a routing loop inside one of the campus ISP's network.||Certain traffic that traversed this ISP would never make the final destination. Some DNS lookups would have also failed.||Campus was able to route around the problem, and the ISP also corrected their internal problem. The cut fiber was restored last night.|
|2017-03-28 (0000)||2017-03-29 (1600)||LSST||NPCF Chilled Water Outage||LSST - Slurm cluster nodes will be offline during the outage. All other LSST systems are expected to remain operational.||No issues. Slurm nodes restarted.|
|2017-03-28 (0000)||2017-03-29 (0230)||Blue Waters||NPCF Chilled Water Outage||Full system shutdown on Blue Waters (except Sonexion which is needed for fsck)||FSCK done on all lustre file systems, XDP piping works done (no leakage found), Software updates (PE, darshan) completed.|
|Blue Waters||BW scratch MDT failover, df hangs||BW scratch MDT failover, load on mds was 500+ delayed failover. Post FO had some issues that delayed RTS.||scheduler was paused|
|Blue Waters||BW login node ps hang||rebooted h1-h3, lost bw/h2ologin DNS record, had neteng recreate the record. Had to rotate login in and out of round-robins until all rebooted. User email sent (2).||Logins nodes rebooted|
DNS round-robin changes
|2017-03-23 (1000)||2017-03-23 (1500)||Nebula||NCSA Nebula Outage||Nebula will take an outage to balance and build a more stable setup for the file system. This will require a pause of all instances, and Horizon being unavailable.||File system online and stable. At this time all blocks were balanced and healed.|
|2017-03-16 (0630)||2017-03-16 (1130)||LSST||LSST monthly maintenance||GPFS filesystems will go offline for entire duration of outages. Some systems may be rebooted, especially those that mount one or more of the GPFS filesystems.|
|Blue Waters||Failure on cabinet c9-7, affecting HSN.||Filesystem hung for several minutes.||Scheduler was paused for 50 minutes. |
Warmswap cabinet c9-7.
Nodes on c9-7 are reserved for further diagnosis.
|2017-03-15 09:00||2017-03-15 12:47||Campus Cluster||UPS work at ACB.Reshuffling electrical drops on 10k controllers, storage IB switches and some servers.||Scheduler will be paused for regular jobs. MWT2 and DES will continue run on their nodes.||UPS work at ACB - incomplete (required additional parts)Redistributing power work done.Scheduler was paused for 3hrs 50 mins.|
|2017-03-10 13:00||2017-03-10 18:00||Campus Cluster||ICCP - We lost 10K controllers due to some type of power disturbance at ACB.||ICCP - Lost all filesystem and its a cluster wide outage.||Recovered missing LUNs and rebooted the cluster. Cluster was back in service at 18:00.|
|2017-03-09 0900||2017-03-09 1500||Roger||ROGER planned PM||batch, hadoop, data transfer services & Ambari||system out for 6hrs, DT services out until 0000|
|2017-03-08 19:41||2017-03-08 22:41||Blue Waters||XDP powered off that served the four cabinets |
(c16-10, c17-10, c18-10, c19-10).
|scheduler paused, four rack power cycled.|
moab required a restart, too many down nodes
and itterations were stuck.
|2017-03-03 1700||2017-03-03 2200||Blue Waters||BW hpss emergency outage to clean|
up db2 database
|ncsa#nearline, stores are failing with cache full||Resolved cache full errors|
|2017-02-28 1200||2017-02-28 1250||Campus Cluster||ICC Resource Manager down||User can't submit new jobs or start new jobs||Remove corrupted job file|
|2017-02-22 1615||2017-02-221815||Nebula||Nebula Gluster Issues||All Nebula instances paused while gluster repaired||Nebula is available.|
|2017-02-11 1900||2017-02-11 2359||NPCF||NPCF Power Hit||BW Lustre was down, xdp heat issues.||RTS 2017-02-11 2359|
|2017-02-15 0800||2017-02-15 1800||Campus Cluster||ICC Scheduled PM||Batch jobs and login nodes access|